OpenStudy (anonymous):

What are the odds against randomly picking a green apple from a fruit bowl containing seven green apples, four red apples, and six blue ones?

OpenStudy (anonymous):

Thanks, but how do I figure out the ones that are against.

OpenStudy (princeharryyy):

Do, it this way probability of occuring of event = 7/17 probability of green not occuring = 10/17 odds against = (10/17)/(7/17) =10/7

OpenStudy (anonymous):

Oh, got it! Thank you!

OpenStudy (princeharryyy):

:) u r welcome.